There is an animal farm where chickens and cows live. All together, there are 80 heads and 224 legs. How many chickens and cows are there on the farm?
Answer by Boreal(15235)

Ch+Co=80
2Ch+4Co=224
-2Ch-2Co=-160
2Co=64
Co=32 cows or 128 legs
Ch=80-Co=48 chickens or 96 legs